User Tools

Site Tools


handover:r1735
r1735

Hobart 12m:

Disk VSN: HOB+0118/32000/1024

Data volume at beginning: 16383.712GB

  • 1700UT Experiment started OK (Jonathan)
  • 0743UT E-remote Control output frozen - telescope appears to be working fine (acquiring new targets at correct UT). Restarting e-remote doesn't fix the problem. (Bryn)
  • 0840UT Restarting field system to fix e-remote control and log monitor, Jamie assisting. Scan 103-0841 missed (Bryn)

Katherine 12m:

Disk VSN: WSRT-039/2000/1024

Data volume at beginning: 0GB

  • 1700UT Experiment started OK (Jonathan)
  • 1449UT Missed first 7 seconds of scan 103-1448 due to “onsource status is slewing”, not happening consistently at the moment (Tiege)
  • 1647UT Missed first 10 seconds of scan 103-1646 due to same error, Yg fine this time.(Tiege)

Yarragadee 12m:

Disk VSN: UAA-0034/2000/1024

Data volume at beginning: 0GB

  • 1700UT Experiment started OK (Jonathan)
  • 1449UT Missed first 7 seconds of scan 103-1448 due to “onsource status is slewing”, not happening consistently at the moment (Tiege)
/home/www/auscope/opswiki/data/pages/handover/r1735.txt · Last modified: 2016/04/12 16:48 by Lucia Plank